OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (thesmartestderpalive):

Are you here?

OpenStudy (thesmartestderpalive):

@fox11897

OpenStudy (anonymous):

yes

OpenStudy (thesmartestderpalive):

Alright Solve both of the equations

OpenStudy (anonymous):

yes please

OpenStudy (thesmartestderpalive):

No, I mean YOU Solve both of them, and give me the sum.

OpenStudy (anonymous):

i mean i dont know the answer to it

OpenStudy (thesmartestderpalive):

Sorry, I was dealing with something else

OpenStudy (anonymous):

ok

Directrix (directrix):

I'll solve the first equation to help you get started. Then, you can do the second one. |5x + 6| = 41 means this: 5x + 6 = 41 OR 5x + 6 = -41 5x = 35 5x = -47 x = 7 x = -47/5

Directrix (directrix):

For |5x + 6| = 41, there are two answers. @fox11897 Your turn to find the answers for |2x + 13| = 28